Find the slope m of the line that passes through the points (3, –8) and (–7, –4).
Answer by checkley77(12844)

(3, –8) and (–7, –4).
Slope=(y2-y1)/(x2-x1)
Slope=(-4+8)/(-7-3)
Slope=4/-10=-2/5=-0.4 ans. for the slope.